OpenStudy (anonymous):

Will Give Fan And Medal!! Part C: Use your answer in part B to determine the axis of symmetry for f(x)?

OpenStudy (anonymous):

y=-16(x-(-1))^2+400 which translates to y = -16(t +1)^2 + 400

OpenStudy (anonymous):

@Opcode @Whitemonsterbunny17 @sourwing

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@nikato @Ashleyisakitty

OpenStudy (nikato):

Ok. So you already have the equation in vertex form y = (x-h)2 + k And I read on this site that the axis of symmetry would be x=h

OpenStudy (anonymous):

Oh okay Thank you very Much :)

OpenStudy (nikato):

If you want to read it, it's somewhere in the middle of the page http://www.mathwarehouse.com/geometry/parabola/axis-of-symmetry.php